Fermi National Laboratory
 

Accelerator Update for Jun 2, 2006 - June 5, 2006

 

Friday June 2

The day shift began with the Main Injector (MI), Recycler, Switchyard (SY), NuMI, MiniBooNE, and the Tevatron (TeV) conducting studies, and with the Antiproton Source in access.

At 9:51 AM, Operations turned off the TeV for transductor work. The work was finished by 10:39 AM.

Operations put the TeV into a dry squeeze at 10:56 AM.

At 12:42 PM, several NuMI devices tripped off (E:HV101 and E:LAM61) on a ground fault indication and took down the NuMI beam permit. Operators contacted experts.

Operations turned off the 400 MeV area at 1:30 PM so that Booster experts could work. They completed their work by 2:17 PM.

Pbar experts began running reverse protons at 3:13 PM.

Operations resumed sending beam to NuMI around 3:15 PM.

At 4:08 PM, NuMI tripped off due to a magnet (on QD113 and QF121) over temperature indication. An expert replaced a blown fuse and everything was reset.

NuMI studies ended at 8:03 PM.

At 11:27 PM, Operations reported that the Tevatron Electron Lens would not turn on. An expert investigated and reported that it was a power supply problem.

Saturday June 3

The midnight shift began with TeV and Pbar experts conducting studies.

At 4:55 AM, the TeV aborted due to a power supply trip (T:SQ). It reset okay.

At 7:06 AM, a MI-50 LCW pump tripped off due to a PLC problem. Operations contacted an expert. MI continues to run without this pump; operators monitored the temperatures.

Pbar studies ended at 8:37 AM.

A water technician reported at 9:32 AM that he had found and replaced a blown fuse. The MI-50 LCW pump is back in operation.

Operations began preparing Pbar for and access at 11:36 AM. They were out of access ten minutes later.

Operations began turning Pbar on at 11:54 AM. Pbar studies will begin when the machine is on.

At 12:53 PM, a Pbar Accumulator quadrupole power supply (A:QT) tripped off for the first of four times in fifty-two minutes. Operations called in an expert. Experts continued with Pbar Debuncher studies.

A power supply expert replaced the ground current circuitry for A:QT at 3:57 PM. It came back on line.

At 5:02 PM, NuMI’s HV101 tripped off on a ground fault indication for the first of three times in thirty minutes. It reset okay, but Operations called in an expert to investigate.

A power supply expert reported at 7:01 PM that he replaced the timing generator for HV101.

At 7:34 PM, MI suffered a vacuum burst. Everything reset okay.

At 9:54 PM, HV101 tripped off. It reset okay.

At 11:25 PM, the TeV suffered a beam abort due to beam positions at sector D0.

Sunday June 4

The midnight shift began with MI, Recycler, SY, Pbar, NuMI, MiniBooNE, and the TeV conducting studies, and with NuMI taking beam.

At 6:42 AM, A:QT tripped off on a ground fault indication. It reset okay. Operations informed an expert.

At 8:44 AM, beam started falling out of Booster at injection. Operations found a temperature problem and had beam back by 9:04 AM. Experts continued their studies.

At 11:27 AM, HV101 tripped off for the first of two times in five minutes. It reset okay.

At 11:38 AM, Operations reported that they were not able to run NuMI at a two-second rep rate without their beam permit tripping off. After talking with experts, Operations set the rep rate for four seconds.

At 3:34 PM, A:QT tripped on a ground fault indication. It reset okay.

At 4:26 PM, the MI beam permit began tripping due to beam fall-out. Operations investigated.

At 5:12 PM, the TeV suffered a quench at sector A0 during an aperture scan. Operations investigated and believed the quench was due to a QPM failure. Operations informed experts who came in and investigated.

Operations contacted MI experts at 6:18 PM to come in and help with beam problems.

The TeV expert found a problem with the sector A1 QPM. Operations began turning on the TeV at 6:39 PM.

Operations and TeV experts brought the TeV up to 980 GeV ramp and then put the machine into a dry squeeze at 8:09 PM.

MI experts established beam at 8:46 PM and began to increase intensity.

Operations reported at 8:46 PM that they had established beam in the TeV.

At 9:03 PM, HV101 tripped off for the first of three times in a little over an hour. It reset okay.

Monday June 5

The midnight shift began with MI, Recycler, SY, Pbar, NuMI, MiniBooNE, and the TeV conducting studies, and with NuMI taking beam.

At 12:35 AM, HV101 tripped off for the first of two times over a six and half hour period. It reset okay.

At 4:46 AM, the NuMI Horn tripped off. It reset okay.

At 5:18 AM, the TeV sector B11 separator sparked.

At 5:44 AM, Operations reported that MI has suffered from a massive orbit shift right after a MECAR trip. Operations investigated and could find nothing obvious wrong. They called in experts. MI resumed taking beam, but suffered with beam position problems.

The Plans for Monday and the rest of the Week

The plan is to continue commissioning. The goals for this week include the following: making antiprotons and transferring them to other machines, establishing beam to MiniBooNE, increasing Booster intensity, possibly replacing a Booster magnet (Q800), and allowing D0 an sixteen-hour access on Tuesday or Wednesday.

Machine Reports

Linac

Okay

Booster

The power amplifier on a RF station (BRF1) needs to be replaced. Experts want two hours to match energies with MI; this will leave little beam for any other machines.

Main Injector

Experts are investigating their tune shift.

Antiproton Source

Experts have swapped the equipment used in the prototype stacktail tank. Experts have found a bad quadrupole shunt that needs replacing. They will try to accumulate antiprotons.

TeV

Experts reported that they have pretty good orbits, but they still need some work. They also need to work on the tunes and coupling. Unfortunately, they found a quadrupole magnet (D16) that had been rolled the wrong way during the shutdown. They are conducting aperture scans. The machine is ready for antiprotons.

Recycler

Experts are ready for antiprotons. They have passed beam with MI. There is an expert here from the company that built the Pelletron to help investigate problems. Experts report that emittances seem to be bigger than normal.

Switchyard 120

Experts tested power supplies. They are ready for beam.

MiniBooNE

Experts are ready for beam.

NuMI

Experts reported that the problem with HV101 was not a ground fault. They have started collecting data.

Collider Detector at Fermilab

CDF experts ran DAQ and trigger testing.

D-Zero

Experts discovered that their noise problem was coming from their muon clock system. They need an sixteen-hour access to repair.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/06/2006    email Fermilab